Elizabeth wants to start selling cookies at a local farmers market. The farmers market charges an initial set up fee of $175, and each cookie will cost $0.25 to make.
Which of the following functions relates to the total cost (C) and the number of cookies made (n)?
A) C=175+0.25n
B) C=0.25n-175
C) C=175-0.25n
D) C=2.5n+175
E) C=0.25n

Answer by math_tutor2020(3835)


Answer: Choice A
C = 175+0.25n

===============================================

Reason:

n = number of cookies made
0.25n = cost of making all of those n cookies at $0.25 each (aka 25 cents each)
0.25n+175 = add on the initial set up fee and we arrive at the total cost

0.25n+175 is the same as 175+0.25n since the order of adding numbers doesn't matter (eg: 2+3 = 3+2)

---------------

An example:
n = 200 cookies are made
0.25n = 0.25*200 = 50 dollars in cost so far, ignoring the $175 fee
0.25n+175 = 50+175 = 225 dollars total cost after adding on that initial $175 fee

In other words,
C = 0.25n+175
C = 0.25*200+175
C = 50+175
C = 225
Making n = 200 cookies will cost C = 225 dollars.